Given the functions ๐‘”(๐‘ฅ) = ๐‘ฅ โˆ’ 5 and ๐‘“(๐‘ฅ) = ๐‘’

๐‘ฅ

Determine (๐‘“ โˆ˜ ๐‘”)(5)


1
Expert's answer
2022-03-18T05:37:03-0400

Substitute "x=5" into the expression for "g(x):" "g(5)=5-5=0". Then, substitute "g(5)" into the expression for "f(x)": "f(g(5))=f(0)=e^0=1". Thus, "(f\circ g)(5)=1".